The sum of ages clayman and ato is 25years.In five years time,the age of ato will be twice the age of clayman now.1.how old is clayman
2.how old is Ato

Answer by fractalier(6550)

Let's call their ages NOW, a and c.
Then we have
a + c = 25
or a = 25 - c
In five years we have
a + 5 = 2c
Now substituting we have
25-c + 5 = 2c
30 - c = 2c
30 = 3c
c = 10
Then
a = 25 - 10 = 15